# Long-Term Fairness Inquiries and Pursuits in Machine Learning: A Survey of Notions, Methods, and Challenges

### v2

@@ 3.2 Cumulative Group Fairness
Another line of work Chi et al. (2021); Wen et al. (2021); Yin et al. (2023) addressed the long-term nature of fairness by formalizing surrogate fairness notions grounded in static fairness notions, mirroring approaches found in other fairness aspects such as intersectionality Gohar & Cheng (2023). These studies define long-term bias as the difference of the cumulative sum of group fairness (e.g., allocation rates) at each time step between different protected groups. These works have operationalized most of these notions within the MDP framework, leveraging RL techniques — a prevalent strategy for addressing long-term fairness dynamics Zhang & Liu (2021). An MDP is defined by a tuple $(S,A,\rho,\mathbf{T},r,\gamma)$, where $\mathcal{S}$ is the state space (e.g., income levels), $\mathcal{A}$ is the action space (accept/reject), $\rho$ is the initial state distribution, $\mathbf{T}:S\times A\rightarrow\Delta(S)$ is the state transition function, $R:S\times A\rightarrow\mathbb{R}$ is the immediate reward function (e.g., bank’s earned profits), and $\gamma$ is the discount factor. The goal is to find a policy $\pi:S\rightarrow\Delta(A)$ that maximizes cumulative reward $\eta(\pi):=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right],\text{ where }s_{0}\sim\rho,\,a_{t}\sim\pi(\cdot|s_{t}),\,s_{t+1}\sim T(\cdot|s_{t},a_{t})$. Chi et al. (2021) propose Return Parity, which stipulates that MDPs from distinct demographic groups, while sharing common state and action spaces, should aim to attain roughly equivalent expected rewards under a policy. Formally:
@@
While these notions are promising for modeling fairness beyond a single time step, Xu et al. (2023) highlights that they fail to capture the variation in temporal importance. To address this, the authors propose the Equal Long-term Benefit Rate (ELBERT) Xu et al. (2023) fairness notion, which considers long-term fairness as the ratio of group supply and group demand. They adapt static fairness notion ratios, e.g., Equal Opportunity, by formulating them as the ratio of group supply (qualified and accepted individuals) to group demand (qualified individuals). Unlike previous adaptations of static notions, they first take the sum of group supply and group demand separately before taking a ratio. They highlight multiple cases (lending, medical treatment allocation) where temporal variations in group supply and demand lead to a false sense of fairness.
@@
Long-Term Benefit Rate (Xu et al., 2023). Given distinct initial distributions $\rho_{0}$ and $\rho_{1}$ of two protected groups, an MDP is said to have achieved a Long-Term Benefit Rate if
@@ 3.3 Meritocratic Fairness
A body of research leverages bandit settings to model long-term fairness. Joseph et al. (2016) initially identified how no-regret algorithms (those avoiding negative feedback) in bandit settings may exhibit increasing unfairness over time. In response, they introduced an individual notion of fairness called Meritocratic Fairness, which posits that it is unjust to preferentially select one individual (e.g., for a loan) over another if the former is less qualified than the latter Joseph et al. (2016), modeled in the contextual bandit setting.
@@
Liu et al. (2017) complement the aforementioned work by Joseph et al. (2018; 2016) by incorporating a smoothness constraint and protecting higher qualifications over lower qualifications in an on-average manner. Meritocratic Fairness assumes that $k$ distinct groups, each with unique mapping from features to expected rewards, are already known. Moreover, it assumes one group member is available and always selected in each round, which is unrealistic(Joseph et al., 2018). To relax this assumption, Joseph et al. (2018) introduces the notion of weakly-meritocratic fairness, which does not assume knowledge of protected group membership.
@@ 3.4 Causality-Motivated Long-Term Fairness Inquiry
Causal modeling of underlying dynamics in long-term settings has been recognized as beneficial for analyzing social implications of automated decision-making (Creager et al., 2020; Zhang et al., 2020). To achieve long-term fairness, understanding the interplay of various processes is crucial (Tang et al., 2023b). Tang et al. (2023a) present a detailed causal model, in terms of a directly acyclic graph (DAG) for two key dynamics in decision-distribution interplay: (1) the decision-making process based on data at each time step and (2) the underlying data generating process influencing future distributions. Rather than enforcing fairness criteria at each time step, they propose setting a goal of achieving long-term fairness and evaluating it alongside.
@@ 3.5 Discussion
Existing long-term fairness notions differ based on the modeling framework used and what aspect (e.g., effort, merit) of fairness they capture dynamically. Cumulative fairness, for instance, extends previous static fairness notions to the long-term context by aggregating fairness measures at each time step using RL. However, recurrent RL policy deployments that enforce static fairness notions at each time step may not be universally suitable for real-world scenarios Henderson et al. (2018), necessitating the development of lightweight techniques. Recourse and effort-based notions, adapted from static settings, are only one-step analyses of long-term fairness applied iteratively to address equality of effort. One limitation is that they measure the anticipated effort in response to algorithmic decisions, neglecting the legal (e.g., affirmative action) and philosophical emphasis on actual historical effort. To bridge this gap, future work could focus on ensuring parity in outcomes based on the effort already put in and developing algorithms and datasets to capture it in long-term settings. Moreover, some notions, like meritocratic and effort fairness, are only applicable in specific applications such as hiring, limiting their use. While Tier Balancing is a step towards a generalized natural notion from the dynamic and long-term perspective through causal dynamic modeling, it’s imperative that such notions find support in effective and efficient learning algorithms. Additionally, current notions often fail to support multiple protected groups, which fails to fully capture the nuances of fairness (Gohar & Cheng, 2023). Incorporating multiple dimensions of intersecting identities is an important challenge to solve. Finally, these notions are strictly limited to tabular data, and developing long-term notions in other domains (e.g., image, video, and language) is a promising direction.
@@ 4 Improving Long-Term Fairness in Dynamic Setting
Several approaches have emerged to address bias in dynamic settings, leveraging interactions with the environment to maintain fairness during learning. Table 1 outlines the representative mitigation methods. Broadly, these approaches are categorized based on whether the decision-making influences the environment and, consequently, future actions. Most mitigation methods consider the dynamics in a multi-step sequential learning framework.
@@ 4.1 Decisions Decoupled from Environment
In this setting, decision-making does not impact the dynamics of the environment, and most of the works assume an online learning setting where the data arrives sequentially and the model learns the dynamics on the go Bechavod et al. (2019). The seminal work of Joseph et al. (2016) proposed a no-regret algorithm, based on chained confidence intervals, for dynamic fairness in the stochastic and contextual bandits problem for the meritocratic fairness notion. They prove that the task of learning a provably fair contextual bandit algorithm can be transformed into a “Knows What It Knows” (KWIK) learning algorithm Li et al. (2008) and propose an efficient learning algorithm. Utilizing the MAB framework, Liu et al. (2017) propose a Thompson-Sampling-based algorithm to achieve calibrated fairness where each arm is selected with the probability equal to that of its reward being the largest while satisfying the smoothness constraint for each round (See Sec 3.3), with a reward distribution in Bernoulli. Wang et al. (2021) further extends their work to propose a more efficient Thompson-Sampling-based algorithm by enforcing merit fairness across rounds instead of for each round and an arbitrary reward function. Heidari & Krause (2018) model the sequential individual fairness in an online learning setting. They propose a Consistently Follow the Leader algorithm that always picks the outcome label closest to the prediction among the labels satisfying the consistency constraint. The labels generated in such schema are proved to follow the optimal hypothesis in bounded time steps. Also targeting individual fairness in online learning, Bechavod et al. (2020) reduces the constrained learning into the standard online batch classification problem. Thus, the common Follow-The-Perturbed-Leader algorithm can be applied oracle-efficiently to achieve sublinear Lagrangian regret. Similarly, Gillen et al. (2018) proposes an algorithm where the learner aims to satisfy a fairness constraint defined using an unknown distance metric in an online linear contextual bandit setting. They assume access to only weak feedback where a regulator can flag unfairness but cannot quantify it across individuals. Finally, Patil et al. (2021) proposes a meta-algorithm, which can be applied to any MAB algorithm (e.g., UCB1) to satisfy a fairness constraint that each arm is pulled at least some pre-specified fraction of the times.
@@
Beyond the bandit setting, Hashimoto et al. (2018) presents a distributionally robust optimization (DRO) procedure to mitigate the disparity amplification over time. The core idea is to constrain the group-wise worst-case loss in each iteration; thus, the data examples with high risk are up-weighted. Specifically, let $P$ denote the distribution of data examples and $\mathcal{B}(P,r)$ denote the chi-squared ball around $P$ of radius $r$, measured by a certain $f$-divergence. The DRO loss parameterized by $\theta$ around all possible perturbation with the maximal radius $r$ around $P$ is ${R}_{\text{DRO}}(\theta;r)=\sup_{Q\in\mathcal{B}(P,r)}\mathbb{E}_{Q}\ell(\theta).$ There are two solutions to the above DRO risk. The first solution assumes the $\chi^{2}$-divergence and solves the dual form of the DRO risk in the following form $\hat{\theta}_{\lambda}=\min_{\theta}\mathbb{E}_{P}[\ell(\theta)-\lambda]_{+}^{2}.$, where $\lambda$ is a dual parameter. The other solution assumes that one can compute the worst case probability distribution $Q^{*}(\theta)\in\operatorname*{arg\,max}_{Q\in\mathcal{B}(P,r)}\mathbb{E}_{Q}\ell(\theta)$ for a given parameter $\theta$ through projected gradient ascent and alternate the optimal $\theta^{*}(Q)\in\operatorname*{arg\,min}_{\theta}\mathbb{E}_{Q}\ell(\theta)$. The final solution $\theta^{*}$ and $Q^{*}$ will be given by the saddle point of the optimization problem. Moreover, the authors demonstrate that utilizing a one-step minimizer during each iteration, reduces the online learning problem into the standard learning problem. Similarly, Hu & Zhang (2022) proposed a repeated risk minimization framework to sequentially mitigate long-term fairness on post-intervention data samples.
@@ 4.2 Decision Impacts Environment Dynamics
A series of works Yin et al. (2023); Chi et al. (2021) model the long-term fairness issue in the context of RL. These works typically focus on how the deployment of the decision policy influences the environment and future rewards. In this context, Jabbari et al. (2017) propose a fair polynomial time algorithm using the MDP framework to enforce the notion of approximate meritocratic fairness. Here, the distribution of individuals is considered as the state $s$ of the environment, and the decision model $f$ is considered as the action $a$. The decision model is evaluated by the risk $\ell(s,a)$ and the fairness violation $G(s,a)$ on the data observation. The utility function for the RL agent is the negative of risk and fairness violation subject to the observed state and the selected decision policy, i.e., $r(s,a)=1-\ell(s,a)$ and $g(s,a)=1-G(s,a)$, respectively. The authors assume an unknown state transition probability such that $s_{t+1}\sim P(s_{t},a_{t})$. In the RL context, the goal of long-term fairness can be viewed as maximizing the cumulative reward $\max_{\pi}\mathbb{E}[\sum_{t=0}^{T}r(s_{t},a_{t})\mid s_{0}]$ while preserving the fairness violation over time $\mathbb{E}[\sum_{t=0}^{T}g(s_{t},a_{t})\mid s_{0}]\geq c.$
Yin et al. (2023) show that a modified variant of the Upper-Confidence Bound algorithm can address the long-term fairness with continuous state and action spaces. In particular, the solution is built on the assumption that the $Q$-function can be parameterized as a linear form with respect to some weight vector $w$ so that $w$ can be precisely solved by a least square problem instead. Wen et al. (2021) proposed a slightly different MDP for modeling long-term fairness. Instead of using a shared reward function for the policymaker, they assumed an additional reward function for individuals. Then, the fairness is measured as the difference of the expected individual reward between groups with the deployed model policy $\pi$. To mitigate fairness, Wen et al. (2021) propose prioritizing the policies closer to holding the fairness constraints and ranking the candidates in terms of their rewards. The sampled policies are then used to further update the decision model’s parameters. The algorithm is proven to be near optimal while enforcing the static fairness constraints. On the other hand, Puranik et al. (2022) propose maximizing applicants’ accumulated qualification scores with policy learning. Lastly, Siddique et al. (2020) enforces the fairness constraint that maximizes the utility of the worst of individuals by formulating it as the generalized Gini social welfare function Weymark (1981). Instead of MDP, they adopt multi-objective MDPs, where rewards take the form of vector instead of scalar, to impose the specific notion of fairness.

### v3

@@ 3.2 Cumulative Group Fairness
Another line of work (Chi et al., 2021; Wen et al., 2021; Yin et al., 2023) addressed the long-term nature of fairness by formalizing surrogate fairness notions grounded in static fairness notions, mirroring approaches found in other fairness aspects such as intersectionality (Gohar & Cheng, 2023). These studies define long-term bias as the difference of the cumulative sum of group fairness (e.g., allocation rates) at each time step between different protected groups. These works have operationalized most of these notions within the MDP framework, leveraging RL techniques — a prevalent strategy for addressing long-term fairness dynamics (Zhang & Liu, 2021). An MDP is defined by a tuple $(S,A,\rho,\mathbf{T},r,\gamma)$, where $\mathcal{S}$ is the state space (e.g., income levels), $\mathcal{A}$ is the action space (accept/reject), $\rho$ is the initial state distribution, $\mathbf{T}:S\times A\rightarrow\Delta(S)$ is the state transition function, $R:S\times A\rightarrow\mathbb{R}$ is the immediate reward function (e.g., bank’s earned profits), and $\gamma$ is the discount factor. The goal is to find a policy $\pi:S\rightarrow\Delta(A)$ that maximizes cumulative reward $\eta(\pi):=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right],\text{ where }s_{0}\sim\rho,\,a_{t}\sim\pi(\cdot|s_{t}),\,s_{t+1}\sim T(\cdot|s_{t},a_{t})$. Chi et al. (2021) propose Return Parity, which stipulates that MDPs from distinct demographic groups, while sharing common state and action spaces, should aim to attain roughly equivalent expected rewards under a policy. Formally:
@@
While these notions are promising for modeling fairness beyond a single time step, Xu et al. (2023b) highlights that they fail to capture the variation in temporal importance. To address this, the authors propose the Equal Long-term Benefit Rate (ELBERT) fairness notion, which considers long-term fairness as the ratio of group supply and group demand (Xu et al., 2023b). They adapt static fairness notion ratios, e.g., Equal Opportunity, by formulating them as the ratio of group supply (qualified and accepted individuals) to group demand (qualified individuals). Unlike previous adaptations of static notions, they first take the sum of group supply and group demand separately before taking a ratio. They highlight multiple cases (lending, medical treatment allocation) where temporal variations in group supply and demand lead to a false sense of fairness.
@@
Long-Term Benefit Rate (Xu et al., 2023b). Given distinct initial distributions $\rho_{0}$ and $\rho_{1}$ of two protected groups, an MDP is said to have achieved a Long-Term Benefit Rate if
@@ 3.3 Meritocratic Fairness
A body of research leverages bandit settings to model long-term fairness. Joseph et al. (2016) initially identified how no-regret algorithms (those avoiding negative feedback) in bandit settings may exhibit increasing unfairness over time. In response, they introduced an individual notion of fairness called Meritocratic Fairness, which posits that it is unjust to preferentially select one individual (e.g., for a loan) over another if the former is less qualified than the latter (Joseph et al., 2016), modeled in the contextual bandit setting.
@@
Liu et al. (2017) complement the aforementioned work by Joseph et al. (2016; 2018) by incorporating a smoothness constraint and protecting higher qualifications over lower qualifications in an on-average manner. Meritocratic Fairness assumes that $k$ distinct groups, each with a unique mapping from features to expected rewards, are already known. Moreover, it assumes one group member is available and always selected in each round, which is unrealistic (Joseph et al., 2018). To relax this assumption, Joseph et al. (2018) introduce the notion of weakly-meritocratic fairness, which does not assume knowledge of protected group membership.
@@ 3.4 Causality-Motivated Long-Term Fairness Inquiry
Causal modeling of underlying dynamics in long-term settings has been recognized as beneficial for analyzing social implications of automated decision-making (Creager et al., 2020; Zhang et al., 2020). To achieve long-term fairness, understanding the interplay of various processes is crucial (Tang et al., 2023b). Tang et al. (2023a) present a detailed causal model, in terms of a directly acyclic graph (DAG) for two key dynamics in decision-distribution interplay: (1) the decision-making process based on data at each time step and (2) the underlying data-generating process influencing future distributions. Rather than enforcing fairness criteria at each time step, they propose setting a goal of achieving long-term fairness and evaluating it alongside.
@@ 3.5 Discussion
Existing long-term fairness notions differ based on the modeling framework used and what aspect (e.g., effort, merit) of fairness they capture dynamically. Cumulative fairness, for instance, extends previous static fairness notions to the long-term context by aggregating fairness measures at each time step using RL. However, recurrent RL policy deployments that enforce static fairness notions at each time step may not be universally suitable for real-world scenarios (Henderson et al., 2018), necessitating the development of lightweight techniques. Recourse and effort-based notions, adapted from static settings, are only one-step analyses of long-term fairness applied iteratively to address equality of effort. One limitation is that they measure the anticipated effort in response to algorithmic decisions, neglecting the legal (e.g., affirmative action) and philosophical emphasis on actual historical effort. To bridge this gap, future work could focus on ensuring parity in outcomes based on the effort already put in and developing algorithms and datasets to capture it in long-term settings. Notably, while effort-based notions consider how individuals respond to incentives, they generally assume straightforward improvements and do not capture complex strategic adaptation, where behavior may change in unexpected or adversarial ways to influence outcomes (Zhang et al., 2022). Incorporating this strategic dimension is an important direction for future research. Moreover, some notions, like meritocratic and effort fairness, are only applicable in specific applications, such as hiring, limiting their use. While Tier Balancing is a step towards a generalized natural notion from the dynamic and long-term perspective through causal dynamic modeling, it’s imperative that such notions find support in effective and efficient learning algorithms. Additionally, current notions often fail to support multiple protected groups, which fails to fully capture the nuances of fairness (Gohar & Cheng, 2023). Incorporating multiple dimensions of intersecting identities is an important challenge to solve. One interesting approach is to extend the Max-Min fairness notion (Ghosh et al., 2021) to long-term settings, optimizing outcomes for the worst-off intersectional subgroups. Finally, these notions are strictly limited to tabular data, and developing long-term fairness notions in other domains (e.g., image, video, and text) is a promising direction.
@@ 4 Improving Long-Term Fairness in Dynamic Setting
Several approaches have emerged to address bias in dynamic settings, leveraging interactions with the environment to maintain fairness during learning. Table 1 outlines the representative mitigation methods. Broadly, these approaches are categorized based on whether the decision-making influences the environment and, consequently, future actions. Most mitigation methods consider the dynamics in a multi-step sequential learning framework. In this setting, the system’s evolution over time is explicitly modeled, with each state depending on past states, actions, and feedback from the environment. This framework captures both immediate and delayed effects, rich feedback loops, and complex interactions, and is the primary setting for mitigation methods.
@@ 4.1 Decisions Decoupled from Environment
In this setting, decision-making does not impact the dynamics of the environment, and most of the works assume an online learning setting where the data arrives sequentially and the model learns the dynamics on the go (Bechavod et al., 2019). The seminal work of Joseph et al. (2016) proposed a no-regret algorithm, based on chained confidence intervals, for dynamic fairness in the stochastic and contextual bandits problem for the meritocratic fairness notion. They prove that the task of learning a provably fair contextual bandit algorithm can be transformed into a “Knows What It Knows” (KWIK) learning algorithm (Li et al., 2008) and propose an efficient learning algorithm. Utilizing the MAB framework, Liu et al. (2017) propose a Thompson-Sampling-based algorithm to achieve calibrated fairness where each arm is selected with the probability equal to that of its reward being the largest while satisfying the smoothness constraint for each round (See Sec 3.3), with a reward distribution in Bernoulli. Wang et al. (2021) further extend their work to propose a more efficient Thompson-Sampling-based algorithm by enforcing merit fairness across rounds instead of for each round and an arbitrary reward function. Heidari & Krause (2018) model the sequential individual fairness in an online learning setting. They propose a Consistently Follow the Leader algorithm that always picks the outcome label closest to the prediction among the labels satisfying the consistency constraint. The labels generated in such a schema are proven to follow the optimal hypothesis in bounded time steps. Also targeting individual fairness in online learning, Bechavod et al. (2020) reduce the constrained learning into the standard online batch classification problem. Thus, the common Follow-The-Perturbed-Leader algorithm can be applied oracle-efficiently to achieve sublinear Lagrangian regret. Similarly, Gillen et al. (2018) propose an algorithm where the learner aims to satisfy a fairness constraint defined using an unknown distance metric in an online linear contextual bandit setting. They assume access to only weak feedback, where a regulator can flag unfairness but cannot quantify it across individuals. Finally, Patil et al. (2021) propose a meta-algorithm, which can be applied to any MAB algorithm (e.g., UCB1) to satisfy a fairness constraint that each arm is pulled at least some pre-specified fraction of the times.
@@
Beyond the bandit setting, Hashimoto et al. (2018) present a distributionally robust optimization (DRO) procedure to mitigate the disparity amplification over time. The core idea is to constrain the group-wise worst-case loss in each iteration; thus, the data examples with high risk are up-weighted. Specifically, let $P$ denote the distribution of data examples and $\mathcal{B}(P,r)$ denote the chi-squared ball around $P$ of radius $r$, measured by a certain $f$-divergence. The DRO loss parameterized by $\theta$ around all possible perturbation with the maximal radius $r$ around $P$ is ${R}_{\text{DRO}}(\theta;r)=\sup_{Q\in\mathcal{B}(P,r)}\mathbb{E}_{Q}\ell(\theta).$ There are two solutions to the above DRO risk. The first solution assumes the $\chi^{2}$-divergence and solves the dual form of the DRO risk in the following form $\hat{\theta}_{\lambda}=\min_{\theta}\mathbb{E}_{P}[\ell(\theta)-\lambda]_{+}^{2}.$, where $\lambda$ is a dual parameter. The other solution assumes that one can compute the worst case probability distribution $Q^{*}(\theta)\in\operatorname*{arg\,max}_{Q\in\mathcal{B}(P,r)}\mathbb{E}_{Q}\ell(\theta)$ for a given parameter $\theta$ through projected gradient ascent and alternate the optimal $\theta^{*}(Q)\in\operatorname*{arg\,min}_{\theta}\mathbb{E}_{Q}\ell(\theta)$. The final solution $\theta^{*}$ and $Q^{*}$ will be given by the saddle point of the optimization problem. Moreover, the authors demonstrate that utilizing a one-step minimizer during each iteration reduces the online learning problem to the standard learning problem. Similarly, Hu & Zhang (2022) propose a repeated risk minimization framework to sequentially mitigate long-term fairness on post-intervention data samples.
@@ 4.2 Decision Impacts Environment Dynamics
A series of works (Yin et al., 2023; Chi et al., 2021) model the long-term fairness issue in the context of RL. These works typically focus on how the deployment of the decision policy influences the environment and future rewards. In this context, Jabbari et al. (2017) propose a fair polynomial time algorithm using the MDP framework to enforce the notion of approximate meritocratic fairness. Here, the distribution of individuals is considered as the state $s$ of the environment, and the decision model $f$ is considered as the action $a$. The decision model is evaluated by the risk $\ell(s,a)$ and the fairness violation $G(s,a)$ on the data observation. The utility function for the RL agent is the negative of risk and fairness violation subject to the observed state and the selected decision policy, i.e., $r(s,a)=1-\ell(s,a)$ and $g(s,a)=1-G(s,a)$, respectively. The authors assume an unknown state transition probability such that $s_{t+1}\sim P(s_{t},a_{t})$. In the RL context, the goal of long-term fairness can be viewed as maximizing the cumulative reward $\max_{\pi}\mathbb{E}[\sum_{t=0}^{T}r(s_{t},a_{t})\mid s_{0}]$ while preserving the fairness violation over time $\mathbb{E}[\sum_{t=0}^{T}g(s_{t},a_{t})\mid s_{0}]\geq c.$
Yin et al. (2023) show that a modified variant of the Upper-Confidence Bound algorithm can address the long-term fairness with continuous state and action spaces. In particular, the solution is built on the assumption that the $Q$-function can be parameterized as a linear form with respect to some weight vector $w$ so that $w$ can be precisely solved by a least squares problem instead. Wen et al. (2021) propose a slightly different MDP for modeling long-term fairness. Instead of using a shared reward function for the policymaker, they assumed an additional reward function for individuals. Then, the fairness is measured as the difference of the expected individual reward between groups with the deployed model policy $\pi$. To mitigate fairness, Wen et al. (2021) propose prioritizing the policies closer to holding the fairness constraints and ranking the candidates in terms of their rewards. The sampled policies are then used to further update the decision model’s parameters. The algorithm is proven to be near optimal while enforcing the static fairness constraints. On the other hand, Puranik et al. (2022) propose maximizing applicants’ accumulated qualification scores with policy learning. Lastly, Siddique et al. (2020) enforces the fairness constraint that maximizes the utility of the worst-off individuals by formulating it as the generalized Gini social welfare function (Weymark, 1981). Instead of MDP, they adopt multi-objective MDPs, where rewards take the form of a vector instead of a scalar, to impose the specific notion of fairness.